3. John Scott has R200 per week available to spend on beer and cigarettes. Beer costs R10 per bottle and cigarettes cost R40 per packet. Suppose John buys 8 bottles of beer and 3 packets of cigarettes and that his marginal utility from beer is then 20 utils and that from cigarettes 35 utils. Is he in equilibrium? If not, should he buy more beer or more cigarettes? Explain.

4. Consider a consumer’s choice between meat and fish. Use indifference curves to illustrate the income effect and substitution effect of an increase in the price of fish.


1
Expert's answer
2020-07-13T15:03:56-0400

The condition of consumer’s equilibrium is

"\frac{Px}{Py}=\frac{MUx}{MUy}"

In this case

"\frac{10}{40}<\frac{20}{35}"

So, John Scott isn’t in equilibrium. He has to buy more beer and less cigarettes, because after that the marginal utility of the last bottle of beer would decrease, while the marginal utility of the last packet of cigarettes would increase, so their ratio would be closer to "\frac{1}{4}"

B)

The allocation of the income effect is achieved by determining its level that would provide the consumer with the opportunity to purchase the same set of goods after changing prices as before the change.

The budget line KL corresponds to cash income I and prices "\text{Px and Py}" . Since fish is presented as product X, we consider the conditions following the increase in the price of product X."\text{
Auxiliary budget direct KL, parallel to KL1, }" not tangential to the previous indifference curve "\text{U1}"

but strictly through the point Е1 corresponding to the optimal set of goods X and Y with the same price ratio.

"\text{The overall result of increasing the price of goods X (X1 - X3) is decomposed into the replacement effect (X1 - X3) and the income effect (X3 - X2)/}"